# Collapsing regular Riemannian foliations with flat leaves

### v1

@@ 3.1. Proof of Theorem B
We fix $p\in M$ and set $k=\dim(\mathcal{F})$. We consider the chart $(W,\phi)$ given by þ2.1 around $p$. For $\overline{p}\in L_{p}\cap W$ observe that
@@
Observe that by construction we have that $\phi^{-1}(U\times\{0,\ldots,0\})=L_{p}\cap W$.
By assuming that $W$ is small enough so that the normal bundle $\nu(M,L_{p})\to L_{p}$ is trivial over $W\cap L_{p}$, we consider over $L_{p}\cap W$ a basis $E_{1},\ldots,E_{m-k}\in\nu(M,L_{p})|_{W\cap L_{p}}$ of smooth vector fields on $L_{p}\cap W=\phi(U\times\{0\})$ normal to $L_{p}$, such that for $\overline{p}\in L_{p}\cap W$ we have $g(\overline{p})(E_{i},E_{j})=\delta_{i,j}$. We take $r>0$ small enough so that $\operatorname{exp}_{\nu}\colon\nu^{r}(M,L_{p})\to M$ is a diffeomorphism onto the open tubular neighborhood $\mathrm{Tub}^{r}(L_{p})$ and set $\overline{W}=W\cap\mathrm{Tub}^{r}(L_{p})$. For $\overline{p}\in L_{p}\cap\overline{W}$ we set $u(\overline{p})=(u_{1}(\overline{p}),\ldots,u_{k}(\overline{p}),0,\ldots,0)$ as a local chart, and observe that for $q\in\overline{W}$ there exists a unique $\overline{p}\in\overline{W}\cap L_{p}$ and unique $t_{1},\ldots,t_{m-k}\in\mathbb{R}$ such that
@@
With these observations we define the Fermi-coordinates $\psi\colon\overline{W}\to\overline{U}\times\overline{V}\subset\mathbb{R}^{k}\times\mathbb{R}^{m-k}$ given as follows: {linenomath}
$\psi_{a}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)$ $=u_{a}(\overline{p})\qquad\mbox{for }a=1,\ldots,k,$
$\psi_{i}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)$ $=t_{i}\qquad\mbox{for }i=k+1,\ldots,m.$
By [24, Lemma 2.4] we have that the restriction of $\frac{\partial}{\partial\psi_{k+1}},\ldots,\frac{\partial}{\partial\psi_{m}}$ to $L_{p}\cap\overline{W}$ are orthonormal. Also we observe that $\frac{\partial}{\partial\psi_{a}}(q)=\frac{\partial}{\partial u_{a}}(q)$, for $a=1,\ldots,k$ by construction.
From now on abusing the notation, we will denote by $x$ the components of $\psi$ of index less than $k$, and by $y$ the components of $\psi$ with index strictly bigger than $k$. That is, $\psi(q)=(x(q),y(q))$ for $q\in\overline{W}$. With this notation we have $\partial/\partial y_{i}=\partial/\partial\psi_{i}$ for $i=k+1,\ldots,m$, and $\partial/\partial x_{a}=\partial/\partial\psi_{a}$ for $a=1,\ldots,k$.
For $(x,y)=q\in\overline{W}$ we have the unique $g$-orthogonal decomposition $\frac{\partial}{\partial y_{i}}=X_{i}+V_{i}$ where $X_{i}\in T_{q}L_{q}$ and $V_{i}\in\nu_{q}(M,L_{q})$. We define matrices $A(x,y)$, $B(x,y)$, $C(x,y)$, and $D(x,y)$ as follows: {linenomath}
$A_{a,b}(x,y)$ $=g(x,y)\left(\frac{\partial}{\partial x_{a}},\frac{\partial}{\partial x_{b}}\right),$
$B_{a,i}(x,y)$ $=g(x,y)\left(\frac{\partial}{\partial x_{a}},\frac{\partial}{\partial y_{i}}\right)=g(x,y)\left(\frac{\partial}{\partial x_{a}},X_{i}\right),$
$C_{i,j}(x,y)$ $=g(x,y)(X_{i},X_{j}),$
$D_{i,j}(x,y)$ $=g(x,y)(V_{i},V_{j}).$
@@
In analogous fashion we have for the metric $g_{\delta}$ and the indices $a,b=1,\ldots,k$, $i,j=k+1,\ldots,m$ the following identities: {linenomath}
@@
Moreover, $A(y)$ induces on $\mathbb{R}^{k}$ a flat metric. Observe that for $r>0$ small enough so that $B_{r}(0)\subset\overline{U}$, and for $\delta$ small enough we have that $g_{\delta}(x,y)=g_{\delta}(y)$ restricted to $B_{r}(0)\times\overline{V}$ is isometric to $g(\bar{x},y)=g(y)$ restricted to $B_{\delta r}(0)\times\overline{V}$. Now by hypothesis
$\lambda\leq\mathrm{Sec}\Big{(}g(x,y)|_{B_{r}(0)\times\overline{V}}\Big{)}\leq\Lambda,$
and thus the same holds for $g(\bar{x},y)|_{B_{\delta r}(0)\times\overline{V}}$. By the compactness of $M$ we get that
@@
Moreover, observe that $\mathrm{Diam}(M,g_{\delta})\leq\mathrm{Diam}(M,g)$, since we are “shrinking some tangent directions” of $M$.
Given a leaf $L_{p}$, we observe that since we are not changing the norm of the tangent directions in $TM$ perpendicular to $L_{p}$, we have that for $r>0$ the tubular neighborhood $B_{r}(L_{p},g)$ of $L_{p}$ with respect to $g$ is the tubular neighborhood $B_{r}(L_{p},g_{\delta})$ of $L_{p}$ with respect to $g_{\delta}$. Moreover we have that $B_{r}(p,g_{\delta})\subset B_{r}(L_{p},g_{\delta})=B_{r}(L_{p},g)$, where $B_{r}(p,g_{\delta})$ is the ball of radius $\delta>0$ centered at $p$ with respect to $g_{\delta}$. But by construction we have that $\mathrm{Vol}(B_{r}(L_{p},g),g_{\delta})$ goes to 00 as $\delta\to 0$. Thus we conclude that
@@
We assume now that there exists $r_{0}>0$ such that $\operatorname{Injrad}(g_{\delta})\geq r_{0}$. By [10, Proposition 14], taking $r_{0}/2>r>0$ we have that there exists a constant $C(m)>0$ depending only on $m$ the dimension of $M$ such that for all $\delta$ we have
$\mathrm{Vol}(B_{r}(p,g_{\delta})\geq C(n)r^{m}>0.$
But this is a contradiction to the fact that $\mathrm{Vol}(B_{r}(p,g_{\delta}))$ decreases to 00. Thus we have that
@@
Thus the metrics $\{g_{\delta}\}_{\delta>0}$ give a collapsing sequence of $M$ with bounded curvature and diameter.
3.2. Proof of Theorem C
In this subsection we present the proof of þC.
Proof of þC.
We take $(M,\mathcal{F},g)$ to be a regular $B$-foliation on a compact $m$-dimensional manifold with finite fundamental group and $m>2$.
We consider the collapsing sequence $\{g_{\delta}\}_{\delta>0}$ given in (3.1) for the proof of þB. Next we fix $p\in M$ and we consider the $(\bar{x},y)$-coordinates on a small neighborhood $\overline{W}$ around $p$ as given in the proof of þB. We recall that in these coordinates we have that
@@ 3.2. Proof of Theorem C
Abusing notation, we now consider the sequence of Riemannian metrics $g_{n}=g_{1/n}$ on $M$, and fix $\varepsilon>0$. Then, by þ2.15 exists an $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ on $M$ with orbits of positive dimension, and there exists $N(\varepsilon)\in\mathbb{N}$ such that for $n\geq N(\varepsilon)$ we have an $\widetilde{\eta}_{\varepsilon}$-invariant Riemannian metric $g_{n}^{\varepsilon}$ on $M$ for which it holds:
@@
We define a Riemannian metric $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ on $M$ as follows: {linenomath}
@@
for $X,Y\in\mathfrak{X}(M)$. For a given $Z\in\mathfrak{X}(M)$ we consider $Z_{\varepsilon}=Z^{\perp}+N(\varepsilon)Z^{\top}$. Then we have that by definition $g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=g(Z,Z)$ and $g^{\varepsilon}_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=\bar{g}^{\varepsilon}_{N(\varepsilon)}(Z,Z)$. Thus from (3.2) we have
$e^{-\varepsilon}g(Z,Z)=e^{-\varepsilon}g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})<g^{\varepsilon}_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})<e^{\varepsilon}g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=e^{\varepsilon}g(Z,Z).$ (3.4)
@@
$e^{-\varepsilon}g(Z,Z)\leq\bar{g}^{\varepsilon}_{N(\varepsilon)}\leq e^{\varepsilon}g(Z,Z).$
@@
This implies that both $\mathrm{Id}\colon(M,d_{g})\to(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ and $\mathrm{Id}^{-1}\colon(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})\to(M,d_{g})$ are $e^{\varepsilon/2}$-Lipschitz, and we have that
$\mathrm{dil}(\operatorname{Id}_{M}),\mathrm{dil}(\operatorname{Id}_{M}^{-1})\leq e^{\varepsilon/2}.$
@@
$d_{L}((M,d_{g}),(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}))\leq e^{\varepsilon/2}.$
This implies that $(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ converges Gromov-Hausdorff to $(M,d_{g})$ as $\varepsilon$ goes to 00.
We recall from þ2.14 that our $\mathsf{N}$-structures $\widetilde{\eta}_{\varepsilon}$ are given by torus fibrations with fiber $T^{k}$ since $M$ has finite fundamental group.
By assuming that the coordinate chart $\overline{W}$ is small enough, we may assume that for the $\mathrm{O}(m)$-principal bundle $\pi\colon F(M)\to M$ we have a trivalization $\psi\colon\pi^{-1}(\overline{W})\to\overline{W}\times\mathrm{O}(m)$. With this trivialization, given $q\in\overline{W}$ we take $\tilde{q}=\psi^{-1}(q,\mathrm{Id})$ to be a lift of $q$, i.e. $\pi(\tilde{q})=q$. We consider the fiber $\widetilde{\eta}_{\varepsilon}^{-1}(\widetilde{\eta}_{\varepsilon}(\tilde{p}))$ of the $\mathsf{N}$-structure that contains $\tilde{q}$, and consider a fixed vector $x\in T_{e}T^{k}$. We now consider the action field $\widetilde{X}_{\varepsilon}(\tilde{q})$ at $\tilde{q}$ of $x$ on the fiber $\widetilde{\eta}^{-1}_{\varepsilon}(\tilde{\eta}_{\varepsilon}(\tilde{q}))$. We set now $X_{\varepsilon}(q)=\pi_{\ast}(\widetilde{X}_{\varepsilon}(\tilde{q}))$; this vector field is well defined, since the fibration $\widetilde{\eta}_{\varepsilon}$ is $\mathrm{O}(m)$-invariant.
@@
Now we observe that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}=\|X_{\varepsilon}^{\perp}\|^{2}_{g_{N(\varepsilon)}}$. Since the orbits of $\widetilde{\eta}_{\varepsilon}$-structure have diameter less than $\varepsilon$ and $x$ is fixed, the norm $\|X_{\varepsilon}\|^{2}_{g^{\varepsilon}_{N(\varepsilon)}}$ goes to 00 as we make $\varepsilon$ tend to 00. From this together with the facts that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}\geq 0$ and $\|X_{\varepsilon}^{\top}\|^{2}_{g_{N(\varepsilon)}}\geq 0$, we conclude that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}\to 0$ as $\varepsilon\to 0$.
In the case when $\|X_{\varepsilon}^{\top}\|^{2}_{g}\to 0$, we conclude that $X_{\varepsilon}$ converges to the origin $\bar{0}\in T_{p}L_{p}$. Thus in this case the tangent space of the $\widetilde{\eta}_{\varepsilon}$-orbits converge to a subspace of the tangent space to the orbits. ∎
To finish the proof of the Theorem C we prove the following lemma.
@@
þ The Riemannian metrics $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ of the previous proof converge uniformly to $g$.
@@
By construction we have that for a given vector $Z\in TM$, the norms $\|Z\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converge to $\|Z\|^{2}_{g}$ as $\varepsilon\to 0$, irrespective of the base point in $M$.
@@
$\bar{g}^{\varepsilon}_{N(\varepsilon)}(X,Y)=\frac{1}{4}\left(\|X+Y\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}-\|X-Y\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}\right),$
@@
we conclude that $\bar{g}^{\varepsilon}_{N(\varepsilon)}(X,Y)\to g(X,Y)$ as $\varepsilon\to 0$ independently at the same rate, independently of the base point $p\in M$.
Consider now a local coordinate system $(x_{1},\ldots,x_{m})$. Then we have that the coefficients $(\bar{g}^{\varepsilon}_{N(\varepsilon)})_{ij}(p)$ converges to $g_{ij}(p)$ with the same rate given by the rate of convergence of $e^{\varepsilon},e^{-\varepsilon}$ to $1$ as $\varepsilon\to 0$, and thus it is independent of the choice of a base point $p$. Thus $(M,\bar{g}^{\varepsilon}_{N(\varepsilon)})$ converges with respect to the $C^{0}$ topology to $(M,g)$. ∎
3.3. Proof of Theorem A
We finish the manuscript by presenting the proof of þA. For this proof we need the following lemma.

### v2

@@ 3.1. Proof of Theorem LABEL:MT:_Collapse_of_flat_foliations
We claim that the family $g_{\delta}$ is the desired collapsing sequence.
We fix $p\in M$ and set $k=\dim(\mathcal{F})$. We consider the chart $(W,\phi)$ given by ?THM? LABEL:P:_Plaque_charts around $p$. For $\overline{p}\in L_{p}\cap W$ observe that
@@
Observe that by construction we have that $\phi^{-1}(U\times\{(0,\ldots,0)\})=L_{p}\cap W$.
By assuming that $W$ is small enough so that the normal bundle $\nu(M,L_{p})\to L_{p}$ is trivial over $W\cap L_{p}$, and thus we can consider a basis $E_{1},\ldots,E_{m-k}\in\nu(M,L_{p})|_{W\cap L_{p}}$ of smooth vector fields on $L_{p}\cap W=\phi(U\times\{\bar{0}\})$ normal to $L_{p}$, such that for $\overline{p}\in L_{p}\cap W$ we have $g(\overline{p})(E_{i},E_{j})=\delta_{i,j}$. We take $r>0$ small enough so that $\operatorname{exp}_{\nu}\colon\nu^{r}(M,L_{p})\to M$ is a diffeomorphism onto the open tubular neighborhood $\mathrm{Tub}^{r}(L_{p})$ and set $\overline{W}=W\cap\mathrm{Tub}^{r}(L_{p})$. For $\overline{p}\in L_{p}\cap\overline{W}$ we set $u(\overline{p})=(u_{1}(\overline{p}),\ldots,u_{k}(\overline{p}),0,\ldots,0)$ as a local chart, and observe that for $q\in\overline{W}$ there exists a unique $\overline{p}\in\overline{W}\cap L_{p}$ and unique $t_{1},\ldots,t_{m-k}\in\mathbb{R}$ such that
@@
With these observations we define the Fermi-coordinates $\psi=(x,y)\colon\overline{W}\to\overline{U}\times\overline{V}\subset\mathbb{R}^{k}\times\mathbb{R}^{m-k}$ with $x(q)=(\psi_{1}(q),\ldots,\psi_{k}(q))$ and $y(q)=(\psi_{k+1}(q),\ldots,\psi_{m}(q))$ given as follows:
$x_{a}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)=\psi_{a}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)$ $=u_{a}(\overline{p})\qquad\mbox{for }a=1,\ldots,k,$
$y_{i}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)=\psi_{i}\left(\operatorname{exp}_{\overline{p}}\left(\sum_{j=1}^{m-k}t_{j}E_{j}(\overline{p})\right)\right)$ $=t_{i}\qquad\mbox{for }i=k+1,\ldots,m.$
By [24, Lemma 2.4] we have that the restriction of $\frac{\partial}{\partial y_{k+1}}=\frac{\partial}{\partial\psi_{k+1}},\ldots,\frac{\partial}{\partial y_{m}}=\frac{\partial}{\partial\psi_{m}}$ to $L_{p}\cap\overline{W}$ are orthonormal. Also we observe that $\frac{\partial}{\partial x_{a}}(q)=\frac{\partial}{\partial\psi_{a}}(q)=\frac{\partial}{\partial u_{a}}(q)$, for $a=1,\ldots,k$ by construction.
For $(x,y)=q\in\overline{W}$ we have the unique $g$-orthogonal decomposition $\frac{\partial}{\partial y_{i}}=X_{i}+V_{i}$ where $X_{i}\in T_{q}L_{q}$ and $V_{i}\in\nu_{q}(M,L_{q})$. We define matrices $A(x,y)$, $B(x,y)$, $C(x,y)$, and $D(x,y)$ as follows:
$A_{a,b}(x,y)$ $=g(x,y)\left(\frac{\partial}{\partial x_{a}},\frac{\partial}{\partial x_{b}}\right),$ $1\leq a,b\leq k,$
$B_{a,i}(x,y)$ $=g(x,y)\left(\frac{\partial}{\partial x_{a}},\frac{\partial}{\partial y_{i}}\right)=g(x,y)\left(\frac{\partial}{\partial x_{a}},X_{i}\right),$ $1\leq a\leq k,\>k+1\leq i\leq m,$
$C_{i,j}(x,y)$ $=g(x,y)(X_{i},X_{j}),$ $k+1\leq i,j\leq m,$
$D_{i,j}(x,y)$ $=g(x,y)(V_{i},V_{j}),$ $k+1\leq i,j\leq m.$
@@
In analogous fashion we have for the metric $g_{\delta}$ and the indices $a,b=1,\ldots,k$, $i,j=k+1,\ldots,m$ the following identities:
@@
Moreover, $A(y)$ induces a flat metric on $\mathbb{R}^{k}$. Observe that for $r>0$ small enough so that $B_{r}(\bar{0})\subset\overline{U}$, and for $\delta$ small enough we have that $g_{\delta}(x,y)=g_{\delta}(y)$ restricted to $B_{r}(\bar{0})\times\overline{V}$ is isometric to $g(\bar{x},y)=g(y)$ restricted to $B_{\delta r}(\bar{0})\times\overline{V}$. Now by hypothesis
$\lambda\leq\mathrm{Sec}\Big(g(x,y)|_{B_{r}(\bar{0})\times\overline{V}}\Big)\leq\Lambda,$
and thus the same holds for $g(\bar{x},y)|_{B_{\delta r}(\bar{0})\times\overline{V}}$. By the compactness of $M$ we get that
@@
Moreover, observe that $\operatorname{Diam}(M,g_{\delta})\leq\operatorname{Diam}(M,g)$, since we are “shrinking some tangent directions” of $M$.
Given a leaf $L_{p}$, we observe that since we are not changing the norm of the tangent directions in $TM$ perpendicular to $L_{p}$, we have that for $r>0$ the tubular neighborhood $B_{r}(L_{p},g)$ of $L_{p}$ with respect to $g$ is the tubular neighborhood $B_{r}(L_{p},g_{\delta})$ of $L_{p}$ with respect to $g_{\delta}$. Moreover we have that $B_{r}(p,g_{\delta})\subset B_{r}(L_{p},g_{\delta})=B_{r}(L_{p},g)$, where $B_{r}(p,g_{\delta})$ is the ball of radius $\delta>0$ centered at $p$ with respect to $g_{\delta}$. But by construction we have that $\mathrm{Vol}(B_{r}(L_{p},g),g_{\delta})$ goes to 0 as $\delta\to 0$. Thus we conclude that
@@
We assume now that there exists $r_{0}>0$ such that $\operatorname{Injrad}(g_{\delta})\geq r_{0}$. By [10, Proposition 14], taking $r_{0}/2>r>0$ we have that there exists a constant $C(m)>0$ depending only on $m$ the dimension of $M$ such that for all $\delta$
$\mathrm{Vol}(B_{r}(p,g_{\delta}))\geq C(n)r^{m}>0.$
But this is a contradiction to the fact that $\mathrm{Vol}(B_{r}(p,g_{\delta}))$ decreases to 0. Thus we have that
@@
Thus the metrics $\{g_{\delta}\}_{\delta>0}$ give a collapsing sequence of $M$ with bounded curvature and diameter. ∎
3.2. Proof of Theorem LABEL:MT:_Approximating_flat_foliations_by_N-structures
In this subsection we present the proof of ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures. The idea of the proof consists in modifying the collapsing sequence from ?THM? LABEL:MT:_Collapse_of_flat_foliations to obtain a new sequence of metrics converging to the original foliated metric $g$. Then we prove that the $g$-perpendicular part of action fields to $\mathcal{F}$ vanishes.
Setup for the proof of Theorem LABEL:MT:_Approximating_flat_foliations_by_N-structures
We give the necessary setup for the proof. We take $(M,\mathcal{F},g)$ to be a regular $B$-foliation on a compact $m$-dimensional manifold with finite fundamental group and $m>2$.
We consider the collapsing sequence $\{g_{\delta}\}_{\delta>0}$ given in (3.1) for the proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations. Next we fix $p\in M$ and we consider the $(\bar{x},y)$-coordinates on a small neighborhood $\overline{W}$ around $p$ as given in the proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations. We recall that in these coordinates we have that
@@ 3.2. Proof of Theorem LABEL:MT:_Approximating_flat_foliations_by_N-structures
Abusing notation, taking $\delta=1/n$ we now consider the sequence of Riemannian metrics $g_{n}=g_{1/n}$ on $M$, and fix $\varepsilon>0$. Then, by ?THM? LABEL:T:_Existence_of_N_structures_on exists an $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ on $M$ with orbits of positive dimension, and there exists $N(\varepsilon)\in\mathbb{N}$ such that for $n\geq N(\varepsilon)$ we have an $\widetilde{\eta}_{\varepsilon}$-invariant Riemannian metric $g_{n}^{\varepsilon}$ on $M$ for which it holds:
@@
We define a Riemannian metric $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ on $M$ as follows:
@@
for $X,Y\in\mathfrak{X}(M)$.
We prove the first part of ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures.
Lemma 3.1.
[…]
Proof.
For a given $Z\in\mathfrak{X}(M)$ we consider $Z_{\varepsilon}=Z^{\perp}+N(\varepsilon)Z^{\top}$. Then we have that by definition $g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=g(Z,Z)$ and $g^{\varepsilon}_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=\bar{g}^{\varepsilon}_{N(\varepsilon)}(Z,Z)$. Thus from (3.2) we have
$e^{-\varepsilon}g(Z,Z)=e^{-\varepsilon}g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})<g^{\varepsilon}_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})<e^{\varepsilon}g_{N(\varepsilon)}(Z_{\varepsilon},Z_{\varepsilon})=e^{\varepsilon}g(Z,Z).$
@@
$e^{-\varepsilon}g(Z,Z)\leq\bar{g}^{\varepsilon}_{N(\varepsilon)}(Z,Z)\leq e^{\varepsilon}g(Z,Z).$ (3.4)
@@
[…]
Recall that the dilatation of a Lipschitz function $f\colon(X,d_{X})\to(Y,d_{Y})$ is $\mathrm{dil}(f)=\sup\{d_{Y}(f(x_{1}),f(x_{2}))/d_{X}(x_{1},x_{2})\mid x_{1},x_{2}\in X\}$. Observe that given two different metrics $d_{1}$, $d_{2}$ on the same space $X$ such that $\operatorname{Id}_{X}\colon(X,d_{1})\to(X,d_{2})$ and $\operatorname{Id}_{X}^{-1}\colon(X,d_{2})\to(X,d_{1})$ are Lipschitz, we have that $\mathrm{dil}(\operatorname{Id}_{X})$ is not necessarily equal to $\mathrm{dil}(\operatorname{Id}_{X}^{-1})$ since this quantities depend on the metrics. Then we have that
$\mathrm{dil}(\operatorname{Id}_{M})\leq e^{\varepsilon/2},\quad\mbox{and}\quad\mathrm{dil}(\operatorname{Id}_{M}^{-1})\leq e^{\varepsilon/2},$
@@
$d_{L}((M,d_{g}),(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}))\leq\varepsilon/2.$
[…]
We recall from ?THM? LABEL:T:_N-structure_over_manifold_with_finite_fundamental_group_are_given_by_T-structure that our $\mathsf{N}$-structures $\widetilde{\eta}_{\varepsilon}$ are given by torus fibrations with fiber $T^{k}$ since $M$ has finite fundamental group.
[…]
We now proof the second part of ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures.
Lemma 3.2.
The vector field $X_{\varepsilon}(q)$ converges to a direction in $T_{q}(L_{q})$.
Proof.
@@
Now we observe that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}=\|X_{\varepsilon}^{\perp}\|^{2}_{g_{N(\varepsilon)}}$. Since the orbits of $\widetilde{\eta}_{\varepsilon}$-structure have diameter less than $\varepsilon$ and $x$ is fixed, the norm $\|X_{\varepsilon}\|^{2}_{g^{\varepsilon}_{N(\varepsilon)}}$ goes to 0 as we make $\varepsilon$ tend to 0. From this together with the facts that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}\geq 0$ and $\|X_{\varepsilon}^{\top}\|^{2}_{g_{N(\varepsilon)}}\geq 0$, we conclude that $\|X_{\varepsilon}^{\perp}\|^{2}_{g}\to 0$ as $\varepsilon\to 0$.
In the case when $\|X_{\varepsilon}^{\top}\|^{2}_{g}\to 0$, we conclude that $X_{\varepsilon}$ converges to the origin $\bar{O}\in T_{p}L_{p}$. Thus in this case the tangent space of the $\widetilde{\eta}_{\varepsilon}$-orbits converge to a subspace of the tangent space to the orbits. ∎
We finish this section by proving that we have also uniform convergence of the Riemannian metrics $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ to $g$.
@@
The Riemannian metrics $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ defined in (3.3) converge uniformly to $g$.
@@
As pointed out in the proof of ?THM? LABEL:L:_modified_metric_converge_GH_to_g, given $Z\in\mathfrak{X}(M)$ we have by (3.4) that the norms $\|Z\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converge to $\|Z\|^{2}_{g}$ as $\varepsilon\to 0$.
@@
[…]
@@
we conclude that $\bar{g}^{\varepsilon}_{N(\varepsilon)}(X,Y)\to g(X,Y)$ as $\varepsilon\to 0$ at the same rate, independently of the base point $p\in M$.
[…]
3.3. Proof of Theorem LABEL:MT:_flat_foliations_are_given_by_group_actions
We finish the manuscript by presenting the proof of ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions. For manifolds of dimension bigger than $2$, the proof consists of showing first that the orbit spaces of the $T^{k}$-actions induced by the $\mathsf{N}$-structures converge in the Gromov-Hausdorff sense to the quotient space $M/\mathcal{F}$. Then by equivariant Gromov-Hausdorff convergence theory for the $T^{k}$-actions there exists a limit action by isometries by a Lie group $G$. Then we prove that the tangent distribution to the $G$-principal orbits agrees with the tangent distribution of $\mathcal{F}$. Lastly, we prove that the orbits of the $G$-action agree with the orbits of the foliation and that $G$ is a torus. The lower dimensional cases are treated separately using existing classification results.
Setup for the Proof of Theorem LABEL:MT:_flat_foliations_are_given_by_group_actions for dimensions >2>2
We present the necessary set up to prove ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions for the case when the dimension of the manifold is >2>2.
Let $(M,\mathcal{F},g)$ be a compact simply-connected Riemannian manifold of dimension >2>2 with a regular $B$-foliation.
In this case the hypothesis of ?THM? LABEL:MT:_Collapse_of_flat_foliations are satisfied, and thus there exists a collapsing sequence of Riemannian metrics $g_{\delta}$ with bounded curvature defined in (3.1). Given $\varepsilon>0$, we consider the $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ given by ?THM? LABEL:T:_Existence_of_N_structures_on. By ?THM? LABEL:T:_N-structure_over_manifold_with_finite_fundamental_group_are_given_by_T-structure it follows that the $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ is given by an effective torus actions $\mu_{\varepsilon}\colon T^{k}\times M\to M$, and we denote by $T^{k}_{\varepsilon}\subset\mathrm{Diff}(M)$ the image of $T^{k}$ under the map $T^{k}\ni\xi\mapsto\mu_{\varepsilon}(\xi,\cdot)\in\mathrm{Diff}(M)$.
[…]
We have the following lemma: